### v3

# MagicDrive3D: Controllable 3D Generation for Any-View Rendering in Street Scenes

Abstract
While controllable generative models for images and videos have achieved remarkable success, high-quality models for 3D scenes, particularly in unbounded scenarios like autonomous driving, remain underdeveloped due to high data acquisition costs. In this paper, we introduce MagicDrive3D, a novel pipeline for controllable 3D street scene generation that supports multi-condition control, including BEV maps, 3D objects, and text descriptions. Unlike previous methods that reconstruct before training the generative models, MagicDrive3D first trains a video generation model and then reconstructs from the generated data. This innovative approach enables easily controllable generation and static scene acquisition, resulting in high-quality scene reconstruction. To address the minor errors in generated content, we propose deformable Gaussian splatting with monocular depth initialization and appearance modeling to manage exposure discrepancies across viewpoints. Validated on the nuScenes dataset, MagicDrive3D generates diverse, high-quality 3D driving scenes that support any-view rendering and enhance downstream tasks like BEV segmentation. Our results demonstrate the framework’s superior performance, showcasing its potential for autonomous driving simulation and beyond.
Figure 1: Rendered panorama of the street scene generated from MagicDrive3D. With conditional controls from 3D bounding boxes of objects, BEV road map, ego trajectory, and text descriptions (e.g., timeofday), MagicDrive3D generates complex open-world 3D scenes represented by deformable Gaussians.
00footnotetext: Corresponding authors.
Project Page: https://flymin.github.io/magicdrive3d/.
1 Introduction
With the advancement of generative models [10, 11, 25, 22], there has been increasing interest in generating 3D assets [19, 28, 2]. While many works focus on object-centric generation [19, 27], generating open-ended 3D scenes remains relatively unexplored. This gap is even more critical considering many downstream applications, such as virtual reality and autonomous driving, require controllable generation of 3D street scenes, which is an open challenge.
The typical workflow for 3D scene generation is geometry-focused [^1] In this paper, we focus on generative models where views/scenes are generated from latent variables.. It begins with reconstructing the scene from multi-view sensors to obtain a compact representation (e.g., voxel grids in NF-LDM [13] and tri-planes in GAUDI [2]), followed by training generative models to produce the scene representation from latent variables. Despite their broad applicability, these methods require expensive data collection for the initial reconstruction step, which sets the upper limit on generation quality. Specifically, they often require static scenes and consistent sensor properties, such as exposure and white balance. Street view datasets, like nuScenes [4], fail to meet these requirements, making it extremely challenging to train geometry-focused 3D street scene generation models using such datasets.
Geometry-free methods [21] represent another category in 3D-aware view synthesis. In street-view generation, multi-view 2D images or videos can be directly generated based on camera parameters [9, 32, 30], excelling in photorealistic image generation. However, they often lack sufficient geometric consistency, limiting their ability to extend to viewpoints beyond the dataset [9, 32, 30].
Recognizing the advancements in controllable generation by geometry-free view synthesis methods [5, 9], it is potential to use them as data engines. Their controllability and photo-realism could address the challenges faced by geometry-focused methods. However, the limited 3D consistency in synthetic views from geometry-free methods, such as temporal inconsistency among frames and deformation of objects, poses crucial issues for integrating two kinds of methods into a unified framework.
To address these challenges, we propose MagicDrive3D, a novel framework that combines geometry-free view synthesis and geometry-focused reconstruction for controllable 3D street scene generation. As illustrated in Figure 2, our approach begins with training a multi-view video generation model to synthesize multiple views of a static scene. This model is configured using controls from object boxes, road maps, text prompts, and camera poses. To enhance inter-frame 3D consistency, we incorporate coordinate embeddings that represent the relative transformation between LiDAR coordinates for accurate control of frame positions. Next, we improve the reconstruction quality of generated views by enhancing 3D Gaussian splatting from the perspectives of prior knowledge, modeling, and loss functions. Given the limited overlap between different camera views [34], we adopt a monocular depth prior and propose a dedicated algorithm for alignment in sparse-view settings. Additionally, we introduce deformable Gaussian splatting and appearance embedding maps to handle local dynamics and exposure discrepancies, respectively.
Demonstrated by extensive experiments, our MagicDrive3D framework excels in generating highly realistic street scenes that align with road maps, 3D bounding boxes, ego trajectory, and text descriptions, as exemplified in Figure 1. We show that the generated camera views can augment training for Bird’s Eye View (BEV) segmentation tasks, improving the view-point robustness on perception tasks. Notably, MagicDrive3D is the first to achieve controllable 3D street scene generation using a common driving dataset (e.g., the nuScenes dataset [4]), without requiring repeated data collection from static scenes.
We summarize our contributions as follows:
  • We propose MagicDrive3D, the first framework to effectively integrate both geometry-free and geometry-focused view synthesis for controllable 3D street scene generation. MagicDrive3D generates realistic 3D street scenes that support rendering from any camera view according to various control signals.
  • We introduce a relative pose embedding technique to generate videos with improved 3D consistency. Additionally, we enhance the reconstruction quality with tailored techniques, including deformable Gaussian splatting, to handle local dynamics and exposure discrepancies in the generated videos.
  • Through extensive experiments, we demonstrate that MagicDrive3D generates high-quality street scenes with multi-dimensional controllability. Our results also show that synthetic data improves 3D perception tasks, highlighting the practical benefits of our approach.
2 Related Work
3D Scene Generation. Numerous 3D-aware generative models can synthesize images with explicit camera pose control [38, 21] and potentially other scene properties [26], but only a few scale for open-ended 3D scene generation. GSN [7] and GAUDI [2], representative of models generating indoor scenes, utilize NeRF [18] with latent code input for “floorplan” or tri-plane feature. Their reliance on datasets covering different camera poses is incompatible with typical driving datasets where camera configuration remains constant. NF-LDM [13] develops a hierarchical latent diffusion model for scene feature voxel grid generation. However, their representation and complex modeling hinder fine detail generation.
Contrary to previous works focusing on scene generation using explicit geometry, often requiring substantial data not suitable for typical street view datasets (e.g., nuScenes [4]) as discussed in Section 1, we propose merging geometry-free view synthesis with geometry-focused scene representations for controllable street scene creation. Methodologically, LucidDreamer [6] is most similar to our approach, although it relies on a text-controlled image generation model, which cannot qualify as a view synthesis model. In contrast, our video generation model is 3D-aware. Besides, we propose several improvements over 3DGS for better scene generation quality. Besides, inpainting-based methods like LucidDreamer [6] and WonderJourney [36] cannot complete our controllable street scene generation task. We showcase the differences in Appendix E.
Street View Video Generation. Diffusion models [25, 11] boost a series of works for street view video generation. These works span from single view video (e.g., [29]) to multi-view video (e.g., [9, 32, 30]). Despite cross-view consistency being essential for multi-view video generation, their generalization ability on camera poses is limited due to the data-centric nature [9]. Furthermore, these models lack exact control over frame transformation (i.e., precise ego trajectory), which is crucial for scene reconstruction. Our work addresses this by enhancing control in video generation and proposing a dedicated deformable Gaussian splatting for geometric assurance.
Street Scene Reconstruction. Scene reconstruction and novel view rendering for street views are useful in applications like driving simulation, data generation, and augmented and virtual reality. For street scenes, attributes like scene dynamic and discrepancies from multi-camera data collection make typical large-scale reconstruction methods ineffective (e.g., [20, 16, 15]). Hence, real data-based reconstruction methods like [34, 35] utilize LiDAR for depth prior, but their output only permits novel view rendering from the same scene. Unlike these methods, our approach synthesizes novel scenes under multiple levels of conditional controls.
3 Preliminaries
Problem Formulation. In this paper, we focus on controllable street scene generation. Given scene description $\mathbf{S}_{t}$, our task is to generate street scenes (represented with 3D Gaussians $\mathbf{G}$) that correspond to the description from a set of latent $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e. $\mathbf{G}=\mathcal{G}(\mathbf{S}_{t},\mathbf{z})$. To describe a street scene, we adopt the most commonly used conditions as per [9, 30, 32]. Specifically, a frame of driving scene $\mathbf{S}_{t}=\{\mathbf{M}_{t},\mathbf{B}_{t},\mathbf{L}_{t}\}$ is described by road map $\mathbf{M}_{t}\in\{0,1\}^{w\times h\times c}$ (a binary map representing a $w\times h$ meter road area in BEV with $c$ semantic classes), 3D bounding boxes $\mathbf{B}_{t}=\{(c_{i},b_{i})\}_{i=1}^{N}$ (each object is described by box $b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$ and class $c_{i}\in\mathcal{C}$), and text $\mathbf{L}_{t}$ describing additional information about the scene (e.g., weather and time of day). In this paper, we parameterize all geometric information according to the LiDAR coordinate of the ego car.
One direct application of scene generation is any-view rendering. Specifically, given any camera pose $\mathbf{P}=[\mathbf{K},\mathbf{R},\mathbf{t}]$ (i.e., intrinsics, rotation, and translation), the model $\mathbf{G}(\cdot)$ should render photo-realistic views with 3D consistency, $\mathcal{I}^{r}=\mathbf{G}(\mathbf{P})$, which is not applicable to previous street view generation (e.g., [9, 30, 32]). Besides, we present more applications in Section 5.
3D Gaussian Splatting. We briefly introduce 3DGS since our scene representation is based on it. 3DGS [12] represents the geometry and appearance via a set of 3D Gaussians $\mathbf{G}$. Each 3D Gaussian is characterized by its position $\boldsymbol{\mu}_{p}$, anisotropic covariance $\boldsymbol{\Sigma}_{p}$, opacity $\alpha_{p}$, and spherical harmonic coefficients for view-dependent colors $c_{p}$. Given a sparse point cloud $\mathcal{P}$ and several camera views $\{\mathcal{I}_{i}\}$ with poses $\{\mathbf{P}_{i}\}$, a point-based volume rendering [41] is applied to make Gaussians optimizable through gradient descend and interleaved point densification. Specifically, the loss is as Equation 1, where $\mathcal{I}^{r}$ is the rendered image, $\lambda$ is a hyper-parameter, and $\mathcal{L}_{\text{D-SSIM}}$ denotes the D-SSIM loss [12].
$\mathcal{L}_{\text{GS}}=(1-\lambda)\mathcal{L}_{1}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})$ (1)
4 Methods
In this section, we introduce our controllable street scene generation pipeline. Due to the challenges that exist in data collection, we integrate geometry-free view synthesis and geometry-focused reconstruction, and propose a generation-reconstruction pipeline, detailed in Section 4.1 and Figure 2. Specifically, we introduce a controllable video generation model to connect control signals with camera views (Section 4.2) and enhance the 3DGS from prior, modeling and loss perspectives (Section 4.3) for better reconstruction with generated views.
Figure 2: Method Overview of MagicDrive3D. For controllable street scene generation, MagicDrive3D decomposes the task into two steps: ① conditional multi-view video generation, which tackles the control signals and provides detailed prior of the scene; and ② scene reconstruction with deformable Gaussian splatting, which guarantees view consistency for any-view rendering.
4.1 3D Street Scene Generation
Direct modeling of controllable street scene generation faces two major challenges: scene dynamics and discrepancy in data collection. Scene dynamics refer to the movements and deformation of elements in the scene, while discrepancy in data collection refer to the discrepancy (e.g., exposure) caused by data collection. These two challenges are even more severe due to the sparsity of cameras for street views (e.g., typically only 6 surrounding perspective cameras). Therefore, reconstruction-generation frameworks do not work well for street scene generation [13, 2].
Figure 2 shows the overview of MagicDrive3D. Given scene descriptions $\mathbf{S}$ as input, MagicDrive3D first extend the descriptions into sequence $\{\mathbf{S}_{t}\}$, where $t\in[0,T]$ according to preset camera poses $\{\mathbf{P}_{c,t}\}$, and generate a sequence of successive multi-view images $\{\mathcal{I}_{c,t}\}$, where $c\in\{1,\dots,N\}$ refers to $N$ surrounding cameras, according to conditions $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$ (detailed in Section 4.2). Then we construct Gaussian representation of the scene with $\{\mathcal{I}_{c,t}\}$ and camera poses $\{\mathbf{P}_{c,t}\}$ as input. This step contains an initializing procedure with a pre-trained monocular depth model and an optimizing process with deformable Gaussian splatting (detailed in Section 4.3). Consequently, the generated street scene not only supports any-view rendering, but also accurately reflects different control signals.
MagicDrive3D integrates geometry-free view synthesis and geometry-focused reconstruction, where control signals are tackled by a multi-view video generator, while the reconstruction step guarantees the generalization ability for any-view rendering. Such a video generator has two advantages: first, since multi-view video generation does not require generalization on novel views [9], it poses less data dependency for street scenes; Second, through conditional training, the model is able to decompose control signals, turning dynamic scenes into static scenes that are more friendly for reconstruction. Besides, for the reconstruction step, strong prior from the multi-view video reduces the burden for scene modeling with complex details.
4.2 Relative Pose Control for Video Generation
Given scene descriptions and a sequence of camera poses $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$, our video generator is responsible for multi-view video generation. Although many previous art for street view generation achieve expressive visual effects, such as [9, 32, 30, 29], their formulations leave out a crucial requirement for 3D modeling. Specifically, the camera pose $\mathbf{P}_{c,t}$ is typically relative to the LiDAR coordinate of each frame. Thus, there is no precise control signal related to the ego trajectory, which significantly determines the geometric relationship between views of different $t$s.
In our video generation model, we amend such precise control ability by adding the transformation between each frame to the first frame, i.e., $\mathbf{T}^{0}_{t}$. To properly encode such information, we adopt Fourier embedding with Multi-Layer Perception (MLP), and concatenate the embedding with the original embedding of $\mathbf{P}_{c,t}$, similar to [9]. As a result, our video generator provides better 3D consistency across frames, most importantly, making the camera poses to each view available in the same coordinate, i.e., $[\mathbf{R}^{0}_{c,t},\mathbf{t}^{0}_{c,t}]=\mathbf{T}^{0}_{t}[\mathbf{R}_{c,t},\mathbf{t}_{c,t}]$.
Figure 3: We optimize the monocular depths (a) with 2 steps for better alignment: coarse scale/offset estimation with SfM PCD (b) and GS optimization (c).
4.3 Enhanced GS for Generated Content
As introduced in Section 3, 3DGS is a flexible explicit representation for 3D reconstruction. Besides, the fast training and rendering speed of 3DGS make it highly suitable for reducing generation costs in our scene creation pipeline. However, similar to other 3D reconstruction methods, 3DGS necessitates high cross-view 3D consistency at the pixel level, which unavoidably magnifies the minute errors in the generated data into conspicuous artifacts. Therefore, we propose improvements for 3DGS from the perspectives of prior, modeling, and loss, enabling 3DGS to tolerate minor artifacts in the generated camera view from our video generation model, thereby becoming a potent tool for enhancing geometric consistency in rendering.
Figure 4: Illustration of the local dynamic from two successive generated frames of Front-Left (FL) camera. Even though our video generation model retains fine 3D consistency, minor discrepancies are inevitable. Our DGS can effectively reconstruct the scene with awareness of such discrepancy.
Prior: Consistent Depth Prior. As essential geometry information, depth is extensively utilized in street scene reconstruction, such as the depth value from LiDAR or other depth sensors used in [34, 35]. However, for synthesized camera views, LiDAR/depth information is unavailable (also making these methods not suitable for our scenario). Therefore, we propose to use pre-trained monocular depth estimator [3] to infer depth information.
While monocular depth estimation is separate for each camera view, proper scale $s_{c,t}$ and offset $b_{c,t}$ parameters should be estimated to align them for a single scene [40], as in Figure 3(a). To this end, we first apply the Point Cloud (PCD) from Structure of Motion (SfM) [24, 23], shown in Figure 3(b). However, such PCD is too sparse to accurately restore $(s_{c,t},b_{c,t})$ for any views. To bridge the final gap, secondly, we propose further optimizing the $(s_{c,t},b_{c,t})$ using the GS loss, as in Figure 3(c). Specifically, we replace the optimization for Guassian centers $\boldsymbol{\mu}_{i}$ with $(s_{c,t},b_{c,t})$. After the optimization, we initialize $\boldsymbol{\mu}_{i}$ with points from depth values. Since GS algorithm is sensitive to accurate point initialization [12, 8], our method provides useful prior to reconstructing in this sparse view scenario.
Modeling: Deformable Gaussians for Local Dynamic. Despite the 3D geometric consistency provided by our video generation model, there are inevitably pixel-level disparities in some object details, as shown in Figure 4. The strict consistency assumption of 3DGS may amplify these minor errors, resulting in floater artifacts. To mitigate the impact of these errors, we propose Deformable Gaussian Splitting (DGS), which, reduces the requirement for temporal consistency between frames, thereby ensuring the reconstruction quality of the generated views.
Specifically, as shown in Figure 4, we pick the center frame $t=t_{C}$ as the canonical space and enforce all Gaussians in this space. Hence, we allocate a set of offsets to each Gaussian, $\boldsymbol{\mu}^{o}_{p}(t)\in\mathbb{R}^{3}$, where $t\in[1,\dots,T]$ and $\boldsymbol{\mu}^{o}_{p}(t_{C})\equiv\mathbf{0}$. Note that, different camera views from the same $t$ share the same $\boldsymbol{\mu}^{o}_{p}(t)$ for each Gaussian, and we apply regularization on them to keep the dynamic in local, as shown in Equation 2:
$\mathcal{L}_{\text{reg}_{o}}=\|\boldsymbol{\mu}^{o}(t)\|_{2}.$ (2)
Consequently, $\boldsymbol{\mu}^{o}_{p}(t)$ can manage the local dynamics driven by pixel-level disparities, while $\boldsymbol{\mu}$ focuses on the global geometric correlations. It ensures the quality of scene reconstruction by leveraging consistent parts across different viewpoints, simultaneously eliminating artifacts. Besides, with the analytical gradient w.r.t. $SE(3)$ pose of cameras [17], we also make the camera pose optimizable in the final few steps of GS iterations, which helps to mitigate the local dynamic from camera poses.
Algorithm 1 Enhanced Deformable GS
1:camera views $\{\mathcal{I}_{i}\}$, camera parameters $\{\mathbf{P}^{0}_{i}\}$, monocular depth $\{\mathcal{D}_{i}\}$, optimization steps for depth $s_{D}$, camera pose $s_{\text{C}}$, and GS $s_{\text{GS}}$
2:DGS of the scene $\{\boldsymbol{\mu}_{p},\boldsymbol{\mu}^{o}_{p},\boldsymbol{\Sigma}_{p},\text{SH}_{p}\}$, and optimized camera pose $\{\mathbf{P}^{0}_{i}\}$
3:$\mathcal{P}_{SfM}=$ PCD from SfM
4:Optimize $(s_{c,t},b_{c,t})$ with $\mathcal{P}_{\text{SfM}}$ for each $\{c,t\}$
5:Random initialize AEs $\{\mathbf{e}_{i}\}$
6:for step in $1,\dots,s_{D}$ do
7:    Random pick one view $\mathcal{I}_{i}$
8:    $\mathcal{L}=\mathcal{L}_{\text{AEGS}}(\mathcal{I}_{i},\mathcal{I}^{r}_{i},\mathbf{e}_{i})$
9:    Update $(s,b),\mathbf{e}_{i},\boldsymbol{\Sigma},\text{SH}$ with $\nabla\mathcal{L}$
10:end for
11:Initialize $\boldsymbol{\mu}$ with $(s,b)$ and $\mathcal{D}$
12:for step in $s_{D},\dots,s_{\text{GS}}$ do
13:    Random pick one view $\mathcal{I}_{i}$ and get its $t$
14:    $\mathcal{L}=\mathcal{L}_{\text{DGS}}(\mathcal{I}_{i},\mathcal{I}^{r}_{i},\mathbf{e}_{i},\boldsymbol{\mu}^{o}(t))$
15:    Update $\boldsymbol{\mu},\boldsymbol{\mu}^{o}(t),\mathbf{e}_{i},\boldsymbol{\Sigma},\text{SH}$ with $\nabla\mathcal{L}$
16:    if step $>s_{\text{C}}$ then
17:         Update $\mathbf{P}^{0}_{i}$ with $\nabla\mathcal{L}$
18:    end if
19:end for
Figure 5: Qualitative comparison with NF-LDM [13] and original 3DGS (on the same generated video as ours). Panoramas for GS are transformed and stitched from perspective cameras with $90^{\circ}$ FOV. Views in the last row are rendered with unseen camera rigs of nuScenes.
Loss: Aligning Exposure with Appearance Modeling. Typical street view datasets are collected with multiple cameras, which capture views independently through auto-exposure and auto-white-balance [4]. Since the video generation is optimized to match the original data distribution, the differences from different cameras also exist in the generated data. The appearance differences are well-known issues for in-the-wild reconstruction [16]. In this paper, we propose a dedicated appearance modeling technique for GS representation.
We hypothesize that the disparity between different views can be represented by affine transformations $\mathbf{A}_{i}(\cdot)$ for $i$-th camera view. An Appearance Embedding (AE) map $\mathbf{e}_{i}\in\mathbb{R}^{w_{e}\times h_{e}\times c_{e}}$ is allocated for each view, and a Convolutional Neural Network (CNN) is utilized to approximate this transformation matrix $w_{\mathbf{A}}\in\mathbb{R}^{w\times h\times 3}$ (Appendix F contains more details). The final computation of the pixel-wise $\ell_{1}$ loss is conducted using the transformed image. Therefore, our final loss for DGS is as Equation 3, where $\lambda_{{\text{reg}_{o}}}$ is the hyper-parameter for offset regularization.
$\begin{split}\mathcal{L}_{\text{DGS}}&=\mathcal{L}_{\text{AEGS}}+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}\\ &=(1-\lambda)\mathcal{L}_{1}(\mathbf{A}_{i}(\mathcal{I}^{r}_{i}),\mathcal{I}_{i})\\ &+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda_{{\text{reg}_{o}}}\mathcal{L}_{{\text{reg}_{o}}}\end{split}$ (3)
Optimization Flow. We demonstrate the overall optimization flow of the proposed DGS in Algorithm 1. Line 2 is the first optimization of monocular depths. Lines 4-8 refer to the second optimization of the monocular depths. Lines 10-16 are the main loop for DGS reconstruction, where we consider temporal offsets on Gaussians, camera pose optimization for local dynamics, and AEs for appearance discrepancies among views.
5 Experiments
5.1 Experimental Setup
Dataset. We test our MagicDrive3D using the nuScenes dataset [4], which is commonly used for generating and reconstructing street views [9, 32, 30, 34]. The official configuration is followed, using 700 street-view video clips of approximately 20s each for training and another 150 clips for validation. For semantics in control signals, we follow [9], using 10 object classes and 8 road classes.
  name #test #train camera poses
  360 6 90 [Uncaptioned image]
vary-t 12 84 [Uncaptioned image]
Table 1: Two settings for reconstruction quality evaluation. Testing views are in green while training views are in red.
Metrics and Settings. MagicDrive3D is primarily evaluated using the Fréchet Inception Distance (FID) by rendering novel views unseen in the dataset and comparing their FID with real images. In addition, the method’s video generation ability is evaluated using Fréchet Video Distance (FVD), and its reconstruction performance is assessed using L1, PSNR, SSIM [31], and LPIPS [37]. For reconstruction evaluation, two testing scenarios are employed: 360, where all six views from $t=9$ are reserved for testing the reconstruction in the canonical space; and vary-t, where one view is randomly sampled from different $t$ to assess long-range reconstruction ability through $t$ in the canonical space (as shown in Table 1). More implementation details can be found in Appendix A.
Methods FVD FID (seen) FID (novel)
MagicDrive[9] 177.26 20.92 N/A
Ours (video gen.) 164.72 20.67 N/A
3DGS N/A 45.07 145.72
Ours (scene gen.) N/A 23.99 34.45
Table 2: Generation quality evaluation. We adopt all validation scenes from nuScenes. All generated views are used for reconstruction. Novel views are camera poses different from nuScenes.
Settings Methods L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
train view vary-t 3DGS 0.0189 30.1191 0.9261 0.1259
3DGS + cc 0.0186 30.2498 0.9253 0.1258
Ours 0.0167 32.6001 0.9544 0.0673
360 3DGS 0.0202 29.4943 0.9187 0.1365
3DGS + cc 0.0199 29.6327 0.9178 0.1366
Ours 0.0174 32.2104 0.9530 0.0693
test view vary-t 3DGS 0.0890 17.9879 0.4378 0.4648
3DGS + cc 0.0799 19.1387 0.4814 0.4697
Ours 0.0738 19.7063 0.5145 0.4115
360 3DGS 0.0910 17.8322 0.4318 0.4756
3DGS + cc 0.0804 19.0773 0.4777 0.4796
Ours 0.0622 21.0351 0.5754 0.3207
Table 3: Reconstruction quality evaluation. We randomly sample 100 scenes from the nuScenes val. set for evaluation. “cc” refers to color correction from [1]. Although 3DGS does not consider appearance differences, we apply “cc” to it for fair comparisons.
5.2 Main Results
Generation Quality. As shown in Table 2, the evaluation of generation quality involves two aspects. Firstly, the quality of video generation is assessed using extended 16-frame MagicDrive [9] as a baseline. Despite minor improvement in single-frame quality based on FID, MagicDrive3D substantially enhances video quality (as evidenced by FVD), demonstrating the efficacy of the proposed relative camera pose embedding in enhancing temporal consistency. Secondly, the image quality of renderings from the generated scene is evaluated using FID. We also include qualitative comparisons in Figure 5. Compared to 3DGS, our enhanced DGS significantly enhances visual quality in reconstructing contents, particularly in unseen novel views. More qualitative comparison can be found in Appendix E.
Reconstruction Quality. Our enhanced DGS, as a reconstruction method, is further evaluated by comparing renderings with ground truth (GT) images. Here, the generated views from the video generator are treated as GT. We employ two settings per Table 1, with results displayed in Table 3. As per all metrics, our enhanced DGS not only improves reconstruction quality for training views but also drastically enhances quality for testing views, compared to 3DGS. We include comparison and discussion with 4DGS [33] in Appendix C.
Methods L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
vary-t setting:
3DGS 0.0799 19.1387 0.4814 0.4697
w/o AE 0.0822 18.8467 0.4758 0.4452
w/o depth scale opt. 0.0815 18.8885 0.4767 0.4366
w/o depth opt. 0.1046 17.2776 0.4399 0.5545
w/o xyz offset + cam 0.0798 19.1657 0.4919 0.4580
Ours 0.0738 19.7063 0.5145 0.4115
360 setting:
3DGS 0.0804 19.0773 0.4777 0.4796
w/o AE 0.0722 19.7742 0.5114 0.3791
w/o depth scale opt. 0.0736 19.6501 0.5086 0.3736
w/o depth opt. 0.0995 17.6707 0.4487 0.5150
w/o xyz offset + cam 0.0798 19.1682 0.4888 0.4663
Ours 0.0622 21.0351 0.5754 0.3207
Table 4: Ablation study on enhanced DGS. We adopt the same settings as in Table 3, where 100 scenes from the nuScenes validation set are adopted.
(a) Original view.
(b) Move forward 0.1m.
(c) Move forward 1m.
(d) Move forward 1.6m.
Figure 6: Application on rendering object-level dynamic. After scene generation, we can segment and move the vehicle (the one on the right) in 3D to render a dynamic object.
Figure 7: Qualitative evaluation for controllability (we show the view from back-left to front-right area). By changing different control signals, MagicDrive3D can edit the scene from different levels.
Controllability. MagicDrive3D accepts 3D bounding boxes, BEV map, and text as control signals, each of which possesses the capacity to independently manipulate the scene. To show such controllability, we edit a scene from the nuScenes validation set, as presented in Figure 7. Clearly, MagicDrive3D can effectively alter the generation of the scene to align with various control signals while maintaining 3D geometric consistency.
Setting Method no rig depth+0.5m pitch-5 yaw+5 yaw-5
vehicle only real data 17.14 16.63 15.50 16.99 15.94
w/ render view (no rig) 20.67 +3.53 20.13 +3.50 17.03 +1.53 19.40 +2.41 19.30 +3.36
w/ random aug. of 4 rigs 21.05 +3.91 20.46 +3.83 19.75 +4.25 19.81 +2.82 19.83 +3.89
road only real data 54.94 54.56 53.82 54.20 53.67
w/ render view (no rig) 60.31 +5.37 59.93 +5.37 58.46 +4.64 59.16 +4.96 59.32 +5.65
w/ random aug. of 4 rigs 60.59 +5.65 60.38 +5.82 59.95 +6.13 60.21 +6.01 60.29 +6.62
Table 5: MagicDrive3D improves the viewpoint robustness [14] of CVT [39]. All results are mIoU for BEV segmentation. Colors highlight the differences with baseline. The best results are in bold.
5.3 Ablation Study
Ablation on Enhanced Gaussian Splatting. As detailed in Section 4.3, three enhancements - prior, modeling, and loss - have been made to 3DGS. To evaluate their efficacy, each was ablated from the final algorithm, the results of which are shown in Table 4. Notations “w/o depth scale opt.” and “w/o depth opt.” represent the absence of GS loss optimization for $(s_{c,t},b_{c,t})$ and use of direct output from the monocular depth model, respectively. Each component’s removal lowered the method’s performance, while incorrect depth sometimes performs worse than the 3DGS baseline. Removal of AE in “vary-t” led to inferior PSNR but improved LPIPS, which is reasonable because AE mitigates the pixel-wise color constraint during reconstruction.
5.4 Application
Training Support for Perception Tasks. We demonstrate an application wherein street scene generation serves as a data engine for perception tasks, leveraging the advantage of any-view rendering ability of our method to improve viewpoint robustness for perception models [14]. We employ CVT [39] and the BEV segmentation task following the evaluation protocols of [39, 9]. By incorporating 4 different rigs on the FRONT camera and adding rendered views for training, the negative impact from viewpoint changes is alleviated (Table 5), exemplifying the utility of street scene generation in training perception tasks.
Render Object-level Dynamic. MagicDrive3D is capable of generating 3DGS representations of scenes, which facilitates various applications in scene editing. Our method utilizes metric scale modeling, guaranteeing that any modifications made to the scenes accurately reflect real-world physical distances. As illustrated in Figure 6, we segmented the generated GS and repositioned the object on the right. The resulting scene GS supports rendering effectively.
6 Conclusion and Discussion
This paper introduces MagicDrive3D, a unique 3D street scene generation framework that integrates geometry-free view synthesis and geometry-focused 3D representations. MagicDrive3D significantly reduces data requirements, enabling training on typical autonomous driving datasets, such as nuScenes. MagicDrive3D employs a video generation model to enhance inter-frame consistency, while the enhanced deformable GS improves reconstruction quality from generated views. Comprehensive experiments demonstrate that MagicDrive3D can produce high-quality 3D street scenes with multi-level controls. Additionally, we show that scene generation can serve as a data engine for perception tasks such as BEV segmentation.
Supplementary Material (Appendix)
Please find the video visualizations on our project website: https://flymin.github.io/magicdrive3d/
A More Implementation Details
For video generation, we train our generator based on the pre-trained street view image generation model from [9]. By adding the proposed relative pose control, we train 4 epochs (77040 steps) on the nuScenes training set with a learning rate of $8e^{-5}$. We follow the settings for 7-frame videos described in [9], using $224\times 400$ for each view but extending to $T=16$ frames. Consequently, for reconstruction, we select $t=8$ as the canonical space. Except we change the first 500 steps to optimize $(s_{c,t},b_{c,t})$ for each view and $\lambda_{\text{reg}_{o}}=1.0$, other settings are the same as 3DGS.
For the monocular depth model, we use ZoeDepth [3]. Although it is trained for metric depth estimation, due to domain differences, raw estimation is not usable, as shown in Section 5.3 and Figure 3. Methodologically, MagicDrive3D does not rely on a specific depth estimation model. Better estimations can further improve our scene generation quality.
Since GS only supports perspective rendering, to stitch the view for panorama, we use code from perspective to equirectangular transformation provided by https://github.com/timy90022/Perspective-and-Equirectangular.
All our experiments are conducted with NVIDIA V100 32GB GPUs. The generation of a single scene takes about 2 minutes for video generation and about 30 minutes for deformable GS reconstruction. For reference, 3DGS reconstruction typically takes about 23 minutes for scenes of similar scales. Therefore, the proposed enhancement is efficient. As for rendering, there is no additional computation for our method compared with 3DGS.
B More Ablation Study
Ablation on Offset Choice for Deformable GS. In addition to the overall module ablation, we observe that for Deformable GS, beyond the Gaussians’ center coordinates, their attributes (including anisotropic covariance, opacity, and harmonic coefficients) can also be utilized to address local inconsistencies. To verify the effects of different choices, we randomly select 10 scenes from the nuScenes validation set for experimentation. As shown in Table I, the results obtained by adding offsets to the center coordinates (xyz) are the best. This aligns with our observation that local inconsistencies in the generated views occur primarily in the shape of objects, and thus, xyz displacements can most effectively resolve these inconsistencies.
Methods L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS 0.0733 19.7514 0.5210 0.4496
Features offset 0.0624 20.9882 0.5940 0.3463
Cov. offset 0.0632 20.8133 0.5854 0.3626
Opacity offset 0.0656 20.5332 0.5733 0.3845
Ours (xyz offset) 0.0546 21.9428 0.6288 0.2759
Table I: Ablating comparison with offsets on anisotropic covariance (Cov.), opacity, and harmonic coefficients (Features) properties in GS. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
C More Reconstruction Baseline
Focusing on dynamic scenes, 4DGS [33] introduces comprehensive improvements over 3DGS and achieves notable results. Therefore, we replace deformable GS with 4DGS. As shown in Table II, by incorporating non-rigid dynamics, 4DGS already performs better than 3DGS. However, in our task, 4DGS underperforms compared to our deformable GS. Based on the results in Table I, we hypothesize that our reconstruction algorithm only needs to address local dynamics caused by content inconsistency. Allocating excessive dynamics at the representational level may hinder model convergence, thus 4DGS does not yield better results in our scenarios.
Methods L1 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS 0.0733 19.7514 0.5210 0.4496
4DGS 0.0601 21.1195 0.5892 0.4475
Ours 0.0546 21.9428 0.6288 0.2759
Table II: Comparison with 4DGS [33]. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
D Limitation and Future Work
As a data-centric method, MagicDrive3D sometimes struggles to generate complex objects like pedestrians, whose appearances are intricate. Additionally, areas with much texture detail (e.g., road fences) or small spatial features (e.g., light poles) are occasionally poorly generated due to limitations in the reconstruction method. Future work may focus on addressing these challenges and further improving the quality and robustness of generated 3D scenes.
E Comparison with Simple Baselines
(a) Conditional image to LucidDreamer [6]
(b) Scene generated by LucidDreamer [6], with text “A driving scene in the city from the front camera of the vehicle. A bus on the right side. There is a bridge overhead. There is a railing in the center of the left road. Some vehicles ahead
(c) Scene generated by WonderJourney [36]. We set the focal length to be the same as the conditional image. WonderJourney cannot control road semantics (many objects are physically implausible) and fails in loop closure for 360 scene generation. Conditions are the same as Figure 1(b).
(d) Stitched panorama with real camera views from nuScenes dataset. Due to the limited overlaps, there are many empty (black) areas.
(e) Panorama from MagicDrive3D. The scene is generated with the same object boxes and BEV map as Figure 1(d), but turned to “rainy day”.
Figure I: Comparison with two baselines (LucidDreamer [6] and WonderJourney [36]) and direct stitching real images.
As shown in Figure I, we further compare MagicDrive3D with two baselines, i.e., LucidDreamer [6] and WonderJourney [36]. The former method has been proposed recently and takes text description as the only condition. Thus, it is hard to generate photo-realistic street scenes. When providing multi-view video frames from nuScenes with known camera poses, their pipeline fails to reconstruct. We suppose the reason is limited overlaps and errors from depth estimation. As suggested by the released code, we changed the image generation model to lllyasviel/control_v11p_sd15_inpaint for inpainting by providing a nuScenes image, i.e., Figure 1(a). However, due to the lack of controllability, the results from LucidDream (e.g., Figure 1(b)) are unsatisfactory. On the other hand, due to the lack of control over objects within the scene, WonderJourney struggles to generate coherent scenes. Inpainting-based methods like the two above exhibit a pronounced sense of patches and face significant challenges in achieving 360 coverage.
Figure 1(d) further shows directly stitching real data. It is also bad due to the limited overlaps between views. On the contrary, the scene generated from MagicDrive3D can render continuous panorama, as shown in Figure 1(e), which is also controllable through multiple conditions.
Note that, panorama generation is only one of the applications of our generated scenes. We show them just for convenient qualitative comparison within the paper. Since our scene generation contains geometric information, they can be rendered from any camera view, as shown in Figure 5.
F Implementation Detail of Appearance Embedding
We show in Figure II the detailed architecture of the CNN used in our appearance modeling. The AE map is 32×32$\times$32 × smaller than the input image to reduce the computational cost. Hence, we first downsample the input image by 32×32$\times$32 ×. Then, we use $3\times 3$ convolution for feature extraction and pixel shuffle for upsampling. Each convolution layer is activated by ReLU.
Figure II: The CNN architecture of appearance modeling, as introduced in Section 4.3.
G Broader Impacts
The implementation of MagicDrive3D in controllable 3D street scene generation could potentially revolutionize the autonomous driving industry. By creating detailed 3D scenarios, self-driving vehicles can be trained more effectively and efficiently for real-world applications, thereby leading to improved safety and accuracy. Moreover, it could potentially provide realistic simulations for human-operated vehicle testing and training, thus contributing to reducing the occurrence of accidents on the roads while enhancing driver expertise. In the broader scope, MagicDrive3D could be of considerable value to the virtual reality industry and video gaming industry, enabling these sectors to generate more lifelike 3D scenes and intricate gaming experiences.
On the downside, the development and application of such advanced technology could lead to certain unwanted scenarios. For instance, the increased automation in industries, driven by the potential of this technology, could lead to job losses for drivers and other related professionals as their roles become automated. A societal transition will be needed to avoid negative impacts on employment levels and the fairness of wealth distribution.
H More Qualitative Results
We show more generated street scenes from MagicDrive3D in Figure III and Figure IV.
Figure III: Generated street scenes from MagicDrive3D. We adopt control signals from the nuScenes validation set. We crop the center part for better visualization.
Figure IV: Generated street scenes from MagicDrive3D. We adopt control signals from nuScenes validation set. The black regions are not fully covered, constrained by the camera’s FOV.